If a^2+6b = _5, b^2 +2c = _6, c^2 +4a = _3 then find a^2+b^2+c^2
 

Grade:10

1 Answers

jagdish singh singh
5 years ago
\hspace{-0.70 cm}a^2+6b+b^2+2c+c^2+4a=-14$\\\\ So $(a^2+4a+4)+(b^2+6b+9)+(c^2+2c+1)=0$\\\\ So $(a+2)^2+(b+3)^2+(c+1)^2=0$ So $a=-2,b=-3,c=-1$\\\\ So $a^2+b^2+c^2=4+9+1=14.$